Is the following expression true or false? Explain your answer.
If r = 9 and s = -4, then r² - 2s = 73.

Thank you!!

Answer by Nate(3500)

If r = 9 and s = -4, then r² - 2s = 73.
Plug the values:
r² - 2s = 73
(9)² - 2(-4) = 73
81 + 8 = 73 False